Are the following lines parallel, perpendicular, or neither?
L1 through (-4,-7) and (1,3)
L2 through (2,6) and (4,10)

Answer by checkley71(8403)

SLOPE FOR L1=(3+7)/(1+4)=10/5=2
SLOPE FOR L2=(10-6)/(4-2)=4/2=2
THESE 2 LINE ARE PARALLEL BECAUSE THEY HAVE THE SAME SLOPE.